# Detecting local perturbations of networks in a latent hyperbolic embedding space

### v1

@@ Abstract
Graph theoretical approaches have been proven to be effective in the characterization of connected systems, as well as in quantifying their dysfunction due to perturbation. In this paper, we show the advantage of a non-Euclidean (hyperbolic) representation of networks to identify local connectivity perturbations and to characterize the induced effects on a large scale. We propose two perturbation scores based on representations of the networks in a latent geometric space, obtained through an embedding onto the hyperbolic Poincaré disk. We numerically demonstrate that these methods are able to localize perturbations in networks with homogeneous or heterogeneous degree connectivity. We apply this framework to identify the most perturbed brain areas in epileptic patients following surgery. This study is conceived in the effort of developing more powerful tools to represent and analyze brain networks, and it is the first to apply geometric network embedding techniques to the case of epilepsy.
@@ 1 Introduction
Any system whose behavior can be seen as resulting from the interaction of a set of interconnected units can be modeled in the mathematical form of a graph. The last two decades have seen a growing interest in applying graph theory to describe the complex structure and function of real-world interconnected systems, such as the Internet, social networks, and the brain. Evidence has revealed[1] that these systems share some peculiar characteristics, like the small-world property [2], hierarchical organization [3], and efficient navigability [4]. Their understanding and analysis would, therefore, require an expansion of standard graph models toward more specific and advanced models and methods.
@@
We can think of the brain as a set of coarse-grained anatomical regions, connected by neuronal pathways, whose density and direction can be mapped comprehensively and non-invasively with neuroimaging techniques, namely magnetic resonance imaging (MRI). The resulting graph encodes this information in the edge weights and represents the brain structure. Mapping brain functions to a graph, instead, involves the recording of brain activities, in the form of blood oxygen level (fMRI) or time-varying electrical or magnetic cortical signals (EEG, MEG). Here, nodes represent the recording sites and the edges encode some measure of correlation between signals. The resulting structural and functional brain networks are the objects of the so-called network neuroscience [5], whose main challenges are the understanding of the emergence of complex behavior and cognition, and the detection and characterization of brain diseases [6].
@@
Indeed, the presence of a neurological disease is usually reflected by alterations of local and global network properties. This has been shown for a wide variety of diseases and disorders like Alzheimer’s disease [7, 8, 9], stroke [10], schizophrenia [11], and epilepsy [12, 13]. Comparing brain networks is a crucial task in many network neuroscience applications [14]. Some diseases can be seen as a network anomaly or perturbation, and the comparison of brain data from a single patient in time (or between groups of diseased and healthy subjects) is the way one can detect and localize the onset of the disease, and the starting point to predict future developments [15]. In recent years for instance, brain network analysis has been applied to improve pre-surgical planning and to predict the post-surgical outcome of patients [13].
@@
Network geometry, particularly hyperbolic geometry, has recently gained attention as a plausible latent geometrical model of complex networks, from which typical properties like a heterogeneous degree distribution and strong clustering can naturally be induced by the metric structure [16]. At the same time, the approach of projecting the network onto a geometric space, called network embedding, has been used to produce highly informative low-dimensional representations of networks. The hypothesis is that complex networks analyses based on these representations can convey additional information on the complex system at a much lower computational cost [17]. When applied to structural brain networks from healthy subjects, embedding in hyperbolic space revealed a multiscale architecture of connectivity and an accurate clustering of anatomical regions in the hyperbolic disk [18]. However, in the case of the study of neurological diseases, this method is still very little used. A recent study on functional brain networks (obtained from magnetoencephalography data) shows its potential in diagnosing early Alzheimer’s states compared to network analysis using connectivity measures alone [19]. However, we are not aware of any studies applying the hyperbolic embedding to epilepsy.
@@
In this work, we study the advantage of non-Euclidean (hyperbolic) representation of networks to identify and characterize local connectivity perturbations. For this, we introduce two methods for comparing a perturbed network with its original version, the details of which are set out in sections §2.2 and §2.2. We assess them on the localization of single-node perturbations on synthetic networks (section §4.1). We finally apply them to localize disrupted areas following an epilepsy surgery (section §4.2).
@@ 2.1 Network embedding in hyperbolic spaces
Embedding networks in Euclidean spaces is a current approach to studying and comparing their connectivity structure [20]. However, such embeddings often require the use of high dimensions and fail to encompass some graph topological properties, such as its hierarchical structure. In contrast, non-Euclidean mapping methods often require a lower number of dimensions to accurately embed graphs and better reflect their organization [21]. Hyperbolic embedding is a non-linear method that takes into account the hierarchical structure of the graph (if any), and depicts both the hierarchies and the similarity of the nodes in the embedding [22]. Embedding in hyperbolic space replicates properties of the graph, such as sparseness, self-similarity, scale-free degree distribution, small-world characteristics, and community structure, that embedding in Euclidean spaces cannot capture [23]. Hyperbolic embeddings have been proven to be useful for community detection and links prediction, as the embedding captures the similarity between network nodes in a graph [24, 23]. Hence, we chose hyperbolic space, rather than Euclidean space, for the network embedding. This decision allowed us to capture the properties of the original network in a lower-dimensional space and quantitatively determine local perturbations to the connectivity structure.
Figure 1: Pipeline for the calculation of hyperbolic perturbation scores. From left to right: network perturbation, mapping of nodes onto the Poincaré disk (lines in gray indicate the links in the original graph), alignment and comparison of original and perturbed network embedding with the HypErr (Method 1) and kNHypD (Method 2) scores. In the right plots, gray curved lines represent the geodesics in the hyperbolic disks.
@@
Hyperbolic geometry studies spaces with a constant negative curvature $K$ that do not conform to Euclidean geometry. When using hyperbolic embedding, the graph is projected onto a hyperboloid that can be further projected onto a 2-dimensional hyperbolic space model like the Poincaré disk or the so-called Klein model disk. Here, we project our connectivity brain networks directly onto the 2D Poincaré disk via the coalescent embedding method [24]. In this algorithm, nodes in the disk are assigned radial coordinates where the radius denotes the centrality of the node, and the angle between them represents their topological proximity.
@@
As a pre-processing step, the weight of edges in the connectivity graph can be adjusted using a repulsion-attraction rule that prioritizes edges with a higher role in information transmission [24]. The resulting matrix is then projected onto a 2D space with a non-linear dimension reduction method, such as Isomap or Laplacian eigenmaps [20]. The nodes are adjusted equally in the polar coordinates by modifying the angular coordinates while maintaining the angular order of the nodes. The final step involves modifying each radius based on its rank, which is given by their degree. Within the hyperbolic disk, the angular distance between two nodes determines their level of similarity, according to a certain definition: for instance, the number of neighbors they have in common. Additionally, each node’s degree centrality is proportional to its distance from the disk’s center: the closer to the center, the more central to the network. For the results presented in this paper, we used the dimensional reduction technique of Isomap, as it is globally optimized and preserves the global network structure by considering the shortest path distances between nodes [20].
@@
Proximity relations in the latent embedding space encode thus the importance of the interaction between nodes. In the case of the brain networks, where nodes correspond to anatomical or functional regions of interest, similar nodes that share, for instance, a high number of topological neighbors, likely belong to the same anatomical cluster and have similar characteristics as computation units in the brain. While nodes with higher centrality usually have a role of integration in the system, interacting with a larger number of anatomical regions. In the hyperbolic embedding, this is encoded in the condition that higher degree nodes are embedded to more central positions in the disk, where the hyperbolic metric naturally ensures that they are typically close, with respect to a more peripheral node, to a larger number of nodes (given that hyperbolic geodesics between peripheral points deviate toward the center).
@@ 2.2 Perturbation detection from network embeddings
$\cosh\mathrm{dist}_{hyp}(i,j)=\cosh r_{i}\times\cosh r_{j}-\sinh r_{i}\times\sinh r_{j}\times\cos(\pi-|\pi-|\theta_{i}-\theta_{j}||)$ (1)
The complete pipeline for scores calculation is shown in Fig. 1.
@@
Method 1: Hyperbolic Error
We define HypErr as the displacement of a node in the hyperbolic disk caused by a perturbation. To estimate this, we embed both the initial and perturbed networks in hyperbolic space using the technique described above in Section § 2.1. Next, we align the two embeddings, as very often, a small perturbation may introduce a random angular offset to the nodes’ positions in the hyperbolic disk. A weak perturbation may thus result in two embeddings with the same similarities between nodes but with a different structure with respect to the coordinates [26]. After aligning both embeddings, a hyperbolic displacement is calculated for each node (see Fig. 1). For each node $i$, its positions resulting from the original and perturbed networks are $Pos_{o}(i)$ and $Pos_{p}(i)$, respectively. The score $\mathrm{HypErr}(i)$ attributed to the node $i$ is then :
$\mathrm{HypErr}(i)=\mathrm{dist}_{hyp}\left(Pos_{o}(i),Pos_{p}(i)\right)$ (2)
Method 2: k-Neighborhood Hyperbolic Distortion
We define the $k$-Neighborhood of node $i$ relative to the hyperbolic embedding of network $G$ as the group of $k$ nodes at a smaller hyperbolic distance to $i$, calling it $kN(i)$. It represents the nodes with which node $i$ has a stronger interaction, as explained in Section §2.1.
To compute the hyperbolic distortion, we consider the same group of nodes in the embedding relative to the perturbed network $G^{\prime}$, denoted as $kN^{\prime}(i)$, and quantify the change in hyperbolic distances due to the perturbation. For the sake of clarity, we will call $i^{\prime}$ the node $i$ in the perturbed network $G^{\prime}$, and we use it to specify that the set of nodes nearest to $i^{\prime}$ in the embedding space of $G^{\prime}$, $kN^{\prime}(i^{\prime})$, is generally different from $kN^{\prime}(i)$. While $kN(i)$ and $kN^{\prime}(i)$ indicate the same network nodes, but in two different embedding configurations, meaning they will generally have different coordinates. These concepts are illustrated on the right side of Fig. 1.
Let’s explain how the method works in the specific: for each node $i$ we compute the set of distances $\{\mathrm{dist}_{hyp}(i,j),j\in kN(i)\}$ in the embedding space of the reference network $G$. We compare these values to $\{\mathrm{dist}_{hyp}(i^{\prime},j),j\in kN^{\prime}(i)\}$ by estimating the score function:
The $\mathrm{kNHypD}(i)$ score reflects the degree of distortion in the $k$-Neighborhood of node $i$ due to a local perturbation of its connectivity. By adding the absolute value, we model as a positive contribution to the score the case in which a distance within the $k$-Neighborhood is reduced in the perturbation. This may occur due to a rewiring process: increasing the degree of a node can push it towards the center of the disk, or increasing the connectivity between two nodes can bring them closer. Dividing the difference by $\mathrm{dist}_{hyp}(i,j)$ gives a higher weight to nodes with which $i$ has a stronger interaction and allows the measure to converge as $k\rightarrow N$. A strong point of this measure is that it is invariant under rotations and reflections of the embedding space; therefore, no prior alignment of the two embeddings to be compared is necessary.
The best choice of $k$ in the estimation of kNHypD depends upon the data. As for other non-parametric estimators in statistics [27], the variance of the estimator appears to be high for small values of $k$. Although increasing the size of the neighborhood reduces the variance, it may also yield very coarse estimates of kNHypD thus increasing the bias. Moreover, the stability gained using higher values of $k$ can be overshadowed by the complexities involved in the computation of a large number of distances. Overall, based on the performances of the estimator across the different networks considered, the size of neighborhood is recommended to be chosen between 10% and 50% of the network size.
@@ 3.1 Synthetic data
We examined the change in the embedding configuration of networks under a “disconnection” perturbation, where nodes are retained (not removed, relabelled, or isolated), but some of their links are cut. This type of disruption mirrors what brain networks experience during epilepsy surgery before any connectivity rewiring mechanism begins. Specifically, we designed a random perturbation method allowing control over the number of target nodes and the size of the perturbation: the percentage of links randomly cut among those available for each target node (see Fig. 2). To ensure method compatibility with real brain networks (see below), we selected synthetic networks with similar size and density: $N=150$ nodes, parameters $k=20$ and $p=0.1$ for Watts-Strogatz, and $m=8$ for the scale-free networks.
Figure 3: Results on synthetic networks. HypErr and kNHypD scores are computed for each node taken as target in a Watts-Strogatz network with $N=150$ nodes, a lattice parameter of $20$ and $0.1$ rewiring probability. We visualise the average values per degree class, measured at the target node itself (in orange), and at the level of its nearest topological neighbors (in green), both relative to the global average over the network, the baseline (dashed grey line). In each subplot $\overline{BL}$ indicates the total average of the scores over all the network’s nodes. The bandwidths indicate the $95\%$ confidence interval of the mean. For the estimation of kNHypD, we set the number of neighbors equal to 10.

### v2

@@ Abstract
Graph theoretical approaches have been proven to be effective in characterizing connected systems, as well as quantifying their dysfunction due to perturbation. In this paper, we show the advantage of a non-Euclidean (hyperbolic) representation of networks to identify local connectivity perturbations and characterize the induced effects on a large scale. We propose two perturbation scores based on representations of the networks in a latent geometric space, obtained through an embedding onto the Poincaré disk. We numerically demonstrate that these scores are able to localize perturbations in networks with homogeneous or heterogeneous degree connectivity. We apply this framework to identify the most perturbed brain areas in epileptic patients following surgery. This study is conceived in the effort of developing more powerful tools to represent and analyze brain networks, and it is the first to apply geometric network embedding techniques to the case of epilepsy.
@@ 1 Introduction
Any system whose behavior can be seen as resulting from the interaction of a set of interconnected units can be modeled in the mathematical form of a graph. The last two decades has seen a growing interest in applying graph theory to describe the complex structure and function of real-world interconnected systems, such as the Internet, social networks, and the brain. Evidence has revealed [1] that these systems share some peculiar characteristics, like the small-world property [2], hierarchical organization [3], and efficient navigability [4]. Therefore, understanding and analyzing them would require an expansion of standard graph models towards more specific and advanced models and methods.
@@
We can think of the brain as a set of coarse-grained anatomical regions, connected by neuronal pathways, whose density and direction can be mapped comprehensively and non-invasively with neuroimaging techniques, namely magnetic resonance imaging (MRI). The resulting graph encodes this information in the edge weights and represents the brain’s structure. Mapping brain functions to a graph, instead, involves the recording of brain activities, in the form of blood oxygen level (fMRI) or time-varying electrical or magnetic cortical signals (EEG, MEG). Here, nodes represent the recording sites and the edges encode some measure of correlation between signals. The resulting structural and functional brain networks are the objects of the so-called network neuroscience [5], whose main challenges are understanding the emergence of complex behavior and cognition and detecting and characterizing brain diseases [6].
@@
Indeed, the presence of a neurological disease is usually reflected by alterations of local and global network properties. This has been shown for a wide variety of diseases and disorders like Alzheimer’s disease [7, 8, 9], stroke [10], schizophrenia [11], and epilepsy [12, 13]. Comparing brain networks is a crucial task in many network neuroscience applications [14]. Some diseases can be seen as a network anomaly or perturbation, and the comparison of brain data from a single patient over time (or between groups of diseased and healthy subjects) is the way one can detect and localize the onset of the disease, and the starting point to predict future developments [15]. In recent years, for instance, brain network analysis has been applied to improve presurgical planning and to predict the postsurgical outcome of patients [13].
Network geometry, particularly hyperbolic geometry, has recently gained attention as a plausible latent geometric model of complex networks, from which typical properties such as a heterogeneous degree distribution and strong clustering can naturally be induced by the metric structure [16]. At the same time, the approach of projecting the network onto a geometric space, called network embedding, has been used to produce highly informative low-dimensional representations of networks. Embedding networks in Euclidean spaces is a current approach to studying and comparing their connectivity structure [17]. However, such embeddings often require the use of high dimensions and fail to encompass some graph topological properties, such as its hierarchical structure. In contrast, non-Euclidean mapping methods often require a lower number of dimensions to accurately embed graphs and better reflect their organization [18].
@@
The hypothesis is that complex network analyses based on these representations can convey additional information on the complex system at a much lower computational cost [19]. When applied to structural brain networks from healthy subjects, embedding in hyperbolic space revealed a multiscale architecture of connectivity and an accurate clustering of anatomical regions in the hyperbolic disk [20]. However, in the case of studying neurological diseases, this method is still rarely used. A recent study on functional brain networks (obtained from magnetoencephalography data) shows its potential in diagnosing early Alzheimer’s states compared to network analysis using connectivity measures alone [21]. However, we are not aware of any studies applying the hyperbolic embedding to epilepsy.
@@
In this work, we study the advantage of non-Euclidean (hyperbolic) representation of networks to identify and characterize local connectivity perturbations. For this, we introduce two scores for comparing a perturbed network with its original version, the details of which are set out in sections §2.2 and §2.2. We assess them on the localization of single-node perturbations on synthetic networks (section §4.1). We finally apply them to localize disrupted areas following an epilepsy surgery (section §4.2).
@@ 2.1 Network embedding in hyperbolic spaces
Hyperbolic embedding is a non-linear method that takes into account the hierarchical structure of the graph (if any), and depicts both the hierarchies and the similarity of the nodes in the embedding [22]. Embedding in hyperbolic space replicates properties of the graph, such as sparseness, self-similarity, scale-free degree distribution, small-world characteristics, and community structure, that embedding in Euclidean spaces cannot capture [23]. Hyperbolic embeddings have been proven to be useful for community detection and link prediction, as the embedding captures the similarity between network nodes in a graph [24, 23]. Therefore, we chose hyperbolic space, instead of Euclidean space, for the network embedding. This decision allowed us to capture the properties of the original network in a lower-dimensional space and quantitatively determine local perturbations to the connectivity structure.
@@
Hyperbolic geometry studies spaces with a constant negative curvature $K$ that do not conform to Euclidean geometry. When using hyperbolic embedding, the graph is projected onto a hyperboloid that can be further projected onto a 2-dimensional hyperbolic space model like the Poincaré disk or the so-called Klein model disk. Current methods to embed networks in a hyperbolic disk essentially belong to two families [25, 22]: generative model-based (e.g. Mercator [26]) and machine learning-based. Here, we project our networks directly onto the 2D Poincaré disk using the coalescent embedding method [24], which belongs to the second family, preferring it for its remarkable versatility and computational speed.
@@
At a glance, the algorithm assigns radial coordinates to the nodes, where the radius denotes the centrality of the node, and the angle between them represents their topological proximity. In a preprocessing step, the weight of edges in the connectivity graph can be adjusted using a repulsion-attraction rule that prioritizes edges with a higher role in information transmission [24]. The resulting matrix is then projected onto a 2D space with a nonlinear dimension reduction method, such as Isomap or Laplacian eigenmaps [17]. The nodes are adjusted equally in the polar coordinates by modifying the angular coordinates while maintaining the angular order of the nodes. The final step involves modifying each radius based on its rank, which is given by their degree. Within the disk, the angular distance between two nodes determines their level of similarity, according to a certain definition: for example, the number of neighbors they have in common. Additionally, each node’s degree centrality is proportional to its distance from the disk’s center: the closer to the center, the more central to the network. An illustration of the hyperbolic embedding of a graph is given in the left frames of Figure 1. For the results presented in this paper, we used the dimensional reduction technique of Isomap, as it is globally optimized and preserves the global network structure by considering the shortest path distances between nodes [17].
@@
Proximity relations in the latent embedding space thus encode the importance of the interaction between nodes. In the case of the brain networks, where nodes correspond to anatomical or functional regions of interest, similar nodes that share, for instance, a high number of topological neighbors, likely belong to the same anatomical cluster and have similar characteristics as computation units in the brain. While nodes with higher centrality usually have a role of integration in the system, interacting with a larger number of anatomical regions. In the hyperbolic embedding, this is encoded in the condition that higher degree nodes are embedded to more central positions in the disk, where the hyperbolic metric naturally ensures that they are typically close, with respect to a more peripheral node, to a larger number of nodes (given that hyperbolic geodesics between peripheral points deviate toward the center).
Figure 1: Pipeline for the calculation of hyperbolic perturbation scores. From left to right: network perturbation, mapping of nodes onto the Poincaré disk (lines in gray indicate the links in the original graph), alignment and comparison of original and perturbed network embedding with the HypDisp (score 1) and kNHypD (score 2) scores. In the right plots, gray curved lines represent the geodesics in the hyperbolic disks.
@@ 2.2 Perturbation detection from network embeddings
$\cosh\mathrm{dist}_{hyp}(i,j)=\cosh r_{i}\times\cosh r_{j}\\ -\sinh r_{i}\times\sinh r_{j}\times\cos(\pi-|\pi-|\theta_{i}-\theta_{j}||)$ (1)
The complete pipeline for scores calculation is shown in Figure 1.
Score 1: Hyperbolic displacement
We define HypDisp as the displacement of a node in the hyperbolic disk caused by a perturbation. To estimate this, we embed both the initial and perturbed networks in hyperbolic space using the technique described above in Section § 2.1. Next, we align the two embeddings, as very often, a small perturbation may introduce a random angular offset to the nodes’ positions in the hyperbolic disk. A weak perturbation may thus result in two embeddings with the same similarities between nodes but with a different structure regarding the coordinates [28]. The alignment is achieved by rotating the perturbed embedding until the minimum sum of HypDisp, calculated between all the nodes of the whole perturbed and original embedded networks, is achieved. After aligning both embeddings, a hyperbolic displacement is calculated for each node (see Figure 1). For each node $i$, its positions resulting from the original and perturbed networks are $Pos_{o}(i)$ and $Pos_{p}(i)$, respectively. The score $\mathrm{HypDisp}(i)$ attributed to the node $i$ is then :
$\mathrm{HypDisp}(i)=\mathrm{dist}_{hyp}\left(Pos_{o}(i),Pos_{p}(i)\right)$ (2)
Score 2: k-Neighborhood Hyperbolic Distortion
We define the $k$-Neighborhood of node $i$ relative to the hyperbolic embedding of network $G$ as the group of $k$ nodes at a smaller hyperbolic distance from $i$, and we call it $kN(i)$. It represents the nodes with which node $i$ has a stronger interaction, as explained in Section §2.1.
To compute the hyperbolic distortion, we consider the same group of nodes in the embedding relative to the perturbed network $G^{\prime}$, denoted as $kN^{\prime}(i)$, and quantify the change in their hyperbolic distances from $i^{\prime}$ due to the perturbation. We call $i^{\prime}$ the node $i$ in the perturbed network $G^{\prime}$, and we use it to specify that the set of nodes nearest to $i^{\prime}$ in the embedding space of $G^{\prime}$, $kN^{\prime}(i^{\prime})$, is generally different from $kN^{\prime}(i)$. We note that $kN(i)$ and $kN^{\prime}(i)$ indicate the same network nodes, but in two different embedding configurations, meaning they will generally have different coordinates. These concepts are illustrated on the right side of Figure 1.
Let’s explain how the score works in the specific: for each node $i$ we compute the set of distances $\{\mathrm{dist}_{hyp}(i,j),j\in kN(i)\}$ in the embedding space of the reference network $G$. We compare these values to $\{\mathrm{dist}_{hyp}(i^{\prime},j),j\in kN^{\prime}(i)\}$ by estimating the score function:
@@
The $\mathrm{kNHypD}(i)$ score reflects the degree of distortion in the $k$-Neighborhood of node $i$ due to a local perturbation of its connectivity. By adding the absolute value, we model a positive contribution to the score in the case where a distance within the $k$-Neighborhood is reduced by the perturbation. This may occur due to a rewiring process: increasing the degree of a node can push it toward the center of the disk, or increasing the connectivity between two nodes can bring them closer. Dividing the difference by $\mathrm{dist}_{hyp}(i,j)$ gives a higher weight to nodes with which $i$ has a stronger interaction. Normalizing by a quadratic factor $\mathrm{dist}^{2}_{hyp}(i,j)$ can ensure convergence of the score for $k\rightarrow N$ and $N\rightarrow\infty$. A strong point of this measure is that it is invariant under rotations and reflections of the embedding space; therefore, no prior alignment of the two embeddings to be compared is necessary.
The best choice of $k$ in the estimation of kNHypD depends upon the data. As for other non-parametric estimators in statistics [29], the variance of the estimator appears to be high for small values of $k$. Although increasing the size of the neighborhood reduces the variance, it may also yield very coarse estimates of kNHypD thus increasing the bias. Moreover, the stability gained by using higher values of $k$ can be overshadowed by the complexities involved in computing a large number of distances. Overall, based on the estimator’s performance across the different networks considered, it is recommended to choose a neighborhood size between 10
@@ 3.1 Synthetic data
We examined the change in the embedding configuration of networks under a “disconnection” perturbation, where nodes are retained (not removed, relabeled, or isolated), but some of their links are cut. This type of disruption mirrors what brain networks experience during epilepsy surgery before any connectivity rewiring mechanism begins. Specifically, we designed a random perturbation method that allows for control over the number of target nodes and the size of the perturbation. This method randomly cuts a percentage of links among those available for each target node(see Figure 2). To ensure method compatibility with real brain networks (see below), we selected synthetic networks with similar size and density: $N=150$ nodes, parameters $k=20$ and $p=0.1$ for Watts-Strogatz, and $m=8$ for the scale-free networks. In our simulations, we excluded realizations with disconnected nodes in order to maintain a connected network.